OpenStudy (anonymous):

according to pauli exclusion principal how we feel electrons in subshails like what is the series 1s 2s 3p what is next

OpenStudy (anonymous):

what?? i didn't get your question.. you are talking about subshells arranging in ascending order of their energies? 1s 2s 2p 3s 3p 4s 3d 4p 5s .. and so on

OpenStudy (anonymous):

ya u r right can you complete this series

OpenStudy (anonymous):

and thanx sir

OpenStudy (anonymous):

the relative energy is computed by adding the main shell and the subshell main shell energy number is given by the main shell number itself subshell energy is given as s=1, p =2, d= 3.. and so on.. so energy of 1s = 1+1 =2.. this is the lowest and comes first then 2s= 2+ 1 = 3 that comes next sometimes the sum comes to be same for ex.. 2p = 2 + 2 and 3s = 3+1.. both give 4.. which has higher energy? the one which has higher main shell has higher energy.. and so on

OpenStudy (anonymous):

k thank u so much

OpenStudy (anonymous):

sure no prob!